OpenStudy (anonymous):

The function p(d)=1+d/33, gives the pressure, in atmosphere (atm), at a depth d in the sea (d is in feet). Note that p(0)=1, p(33)=2, and so on. Find the pressure at 90 feet. The pressure at 90 feet is ___atm. (Type an integer or a simplified fraction).

OpenStudy (anonymous):

Hi jim...i got my answer in decimal, but i need to know how to figure out the simplified fraction...

jimthompson5910 (jim_thompson5910):

p(d)=1+d/33 p(90)=1+90/33 p(90)=33/33+90/33 p(90)=(33+90)/33 p(90)=123/33 So the pressure at 90 ft is 123/33 atm

OpenStudy (anonymous):

so i don't need to divide 123/33 to get the atm?

jimthompson5910 (jim_thompson5910):

oh wait...lol forgot to reduce

jimthompson5910 (jim_thompson5910):

you can divide each term by 3 to get 41/11

jimthompson5910 (jim_thompson5910):

So the answer is 41/11

OpenStudy (anonymous):

haha...divided 123/33 and got 4.72...

jimthompson5910 (jim_thompson5910):

more like 3.72, but they don't want a decimal answer, they want a fraction or an integer

OpenStudy (anonymous):

oh ok, so i had my answer right the first time simplifying

jimthompson5910 (jim_thompson5910):

yeah if you got 41/11

OpenStudy (anonymous):

thats right..yes i did..thank you.

jimthompson5910 (jim_thompson5910):

yw